There is block os m placed on a wedge with slant theta degrees.
Coefficient of friction u.
Find the minimum and maximum value of the horizontal acceleration a of the system so that the block remains stationary with respect to wedge.
 

Grade:11

1 Answers

Arun Kumar IIT Delhi
9 years ago
Hello Student,
\\ \\$forces acting along the surface$ \\mgsin\theta \pm f= ma cos\theta \\f=\mu mgcos\theta \\=>a={mgsin\theta \pm \mu mgcos\theta \over ma cos\theta} \\=>a_{max}={mgsin\theta + \mu mgcos\theta \over ma cos\theta} \\=>a={mgsin\theta - \mu mgcos\theta \over ma cos\theta} \\
Thanks & Regards
Arun Kumar
Btech, IIT Delhi
Askiitians Faculty